OpenStudy (anonymous):

2x - 6y = 5 x + y = 2 Solve the system of equations

OpenStudy (surry99):

rearrange the second equation: x = 2-y Now sub this into the first equation for x... 2(2-y) - 6y = 5 Solve for y, then solve for x

OpenStudy (anonymous):

x = 2-y Now sub this into the first equation for x... 2(2-y) - 6y = 5 4 - 2y - 6y = 5 -8y = 1 y = - 1/8 put in eq.2 x = 2 - (-1/8) x = (16 + 1)/8 x = 17/8